Question 7:

Joe throws a football out into a field for his friend to run and catch. If Joe throws a football at a 30 degree angle with an initial velocity of 10 m/s, how far does his friend have to run in order to catch the football at the same height as Joe throws it and if his friend starts at the same location as he does? Also, how high does the football go relative to the height that Joe throws the football?
Answer #1

Here,

for the same friend to run in order to catch the ball

the horizontal speed of Joe and the football must be same

Horizontal speed of Joe = u * cos(theta)

Horizontal speed of Joe = 10 * cos(30 degree)

Horizontal speed of Joe = 8.66 m/s

=============================

height of the football = (u * sin(theta))^2/(2g)

height of the football = (10 * sin(30 degree))^2/(2 * 9.8)

height of the football = 1.28 m

the height of the football is 1.28 m